# HyperAIRI: a plug-and-play algorithm for precise hyperspectral image reconstruction in radio interferometry

### v1

@@ 5.1 Test dataset
The considered ground-truth images are derived from four radio images of size $N=512\times 512$, which consist of giant radio galaxies 3C 353 (Swain et al., 1998) and Messier 106 (Shimwell et al., 2022), and the radio galaxy clusters Abell 2034 and PSZ2 G165.68+44.01 (Botteon et al., 2022) shown in Fig. 3. The images are first preprocessed to reduce noise and artefacts, then randomly exponentiated to make the final dynamic range in the interval $10^{3},10^{5}]$. We follow the same preprocessing strategies for OAID and MRID and create 40 HDR monochromatic images. Following Thouvenin et al. (2023a), for each image, a spectral index map and a curvature map are generated, in a manner consistent with the procedure used for the training dataset. Specifically, the spectral index map is modelled as the sum of a blurred version of the ground truth in logarithmic scale and a Gaussian field, clipped to the range [5,0][-5,0]. The curvature map is modelled as a Gaussian field, clipped to the range [0.5,0.5][-0.5,0.5]. Each image cube contains $L=36$ channels, with the first channel fixed at the observation frequency of 1 GHz and the last channel randomly chosen from the interval $[1.5,2.0]$ GHz. The observation frequencies of the remaining channels are uniformly spaced within this range. We simulate 40 ground-truth image cubes of size $512\times 512\times 36$ following the spectral model (1).
@@
Hyperspectral RI data are simulated using the antenna configurations of the VLA. The telescope is known for its unique ability to change the positions of the antennas along rail tracks, forming four standard configurations with maximum baselines ranging from 36.4 km to 1.03 km. Configurations with larger baselines provide higher angular resolution, while those with shorter baselines offer higher sensitivity to spatially extended emission. Astronomers often combine measurements from different configurations to obtain a comprehensive Fourier sampling of the target object. Under these considerations, we simulate RI observations combining configurations A and C of the VLA. The Fourier sampling patterns for both configurations are generated using the MeqTrees software package (Noordam & Smirnov, 2010). We randomly vary the telescope’s pointing direction, and the observation duration within the range $[5,10]$ hours, and fix the integration time to 12 seconds. The resulting total number of points in each sampling pattern, at any given spectral channel, varies from $1.1\times 10^{6}$ to $2.1\times 10^{6}$. Examples of simulated sampling patterns for spectral channels with different observation frequencies are shown in Fig. 1. The hyperspectral RI data are simulated under the assumption of a narrow field of view, and following (4), where the standard deviation of the observation noise is fixed across the spectral channels aligning with the target dynamic range of the cube as per the heuristic (14), such that the target dynamic range of the reconstructed image cube matches that of the corresponding ground-truth image cube. The observation’s nominal resolution is parametrised by the super-resolution factor, defined as the ratio between the image spatial Fourier bandwidth and the maximum projected baseline. We fix the super-resolution factor at the highest-frequency channel to 1.5, leading to a super-resolution factor across different channels in the range $[1.5,3.0]$. The resulting test dataset is composed of 40 inverse problems.
Algorithm Joint SNR logSNR RDR sSNR Iterations Resources ${t}_{\text{total}}$ ${t}_{\text{gradient}}$ ${t}_{\text{denoising}}$
deconv. $\pm$ std (dB) $\pm$ std (dB) $\pm$ std $(\times 10^{-3})$ $\pm$ std (dB) $\pm$ std CPUs GPUs $\pm$ std (hr) $\pm$ std (s) $\pm$ std (s)
uSARA No 28.32 $\pm$ 3.24 21.17 $\pm$ 3.56 2.27 $\pm$ 1.32 9.54 $\pm$ 2.34 1251 $\pm$ 299 36 0 2.60 $\pm$ 0.55 2.10 $\pm$ 0.27 5.24 $\pm$ 6.73
AIRI No 28.70 $\pm$ 2.36 20.67 $\pm$ 5.01 2.62 $\pm$ 1.41 9.18 $\pm$ 3.90 3000 $\pm$ 0 36 4 1.84 $\pm$ 0.14 1.60 $\pm$ 0.23 0.46 $\pm$ 0.32
WSClean Yes 11.27 $\pm$ 1.57 12.05 $\pm$ 4.08 2.36 $\pm$ 1.27 3.81 $\pm$ 5.91 10 $\pm$ 1 36 0 0.56 $\pm$ 0.24
Hyper-uSARA Yes 29.30 $\pm$ 2.91 22.30 $\pm$ 4.15 2.32 $\pm$ 1.34 11.35 $\pm$ 3.78 1827 $\pm$ 675 36 0 2.42 $\pm$ 0.87 2.09 $\pm$ 0.27 2.58 $\pm$ 1.43
HyperAIRI Yes 30.53 $\pm$ 1.87 23.07 $\pm$ 5.13 2.44 $\pm$ 1.36 10.87 $\pm$ 3.91 3000 $\pm$ 0 36 4 1.90 $\pm$ 0.14 1.61 $\pm$ 0.23 0.52 $\pm$ 0.25
5.2 Benchmark settings
@@
For benchmarking, we compare HyperAIRI’s performance against four algorithms. These are its optimisation-counterpart Hyper-uSARA, the hyperspectral CLEAN variant in WSClean (Offringa & Smirnov, 2017), and the monochromatic algorithms AIRI and uSARA which are applied for each channel independently. Briggs weighting with robustness parameter set to 0 is considered for all algorithms.
@@
In our experiments, the regularisation parameters for the optimisation algorithms and the effective noise levels for PnP algorithms are directly linked to the heuristic noise level via (14) without further tuning. WSClean parameters are tuned on a small subset of the test dataset and then fixed across the entire dataset to ensure the best overall image quality. In particular, the auto-masking and auto-thresholding values are set to 1.0 and 0.5, respectively, and the number of polynomial terms for spectrum fitting is fixed at 4. The full WSClean command is shown in Appendix B. We also note that the WSClean reconstructions are normalised by the area of the CLEAN beam $A_{\mathrm{beam}}$ on a per-channel basis, thus converting the unit of pixel intensities from Jy/beam to Jy/pixel to align with the groundtruth image cube. The normalisation factor is given by
@@
HyperAIRI, Hyper-uSARA, AIRI and uSARA are implemented in MATLAB, while WSClean is implemented in C++. All experiments were run on Cirrus, a UK Tier 2 HPC service that provides access to both CPU and GPU compute nodes. Each CPU node comprises two Intel 18-core Xeon E5-2695 processors with 256 GB of memory, while each GPU node comprises two 20-core Intel Xeon Gold 6148 processors, four NVIDIA Tesla V100-SXM2-16GB GPUs, and 384 GB of memory. For each inverse problem in the test dataset, uSARA, WSClean and Hyper-uSARA run on a single CPU node, whereas AIRI and HyperAIRI were executed on a single GPU node to accelerate the inference of DNNs. Each AIRI and HyperAIRI denoisers is trained with half GPU node.
5.3 Evaluation metrics
@@
To evaluate the algorithms’ performance in reconstructing faint emissions, we evaluate the SNR on logarithmic scale images, denoted as logSNR. Given a ground-truth image cube, with a dynamic range $a$, we apply the logarithmic scale transform parametrised by $a$ such that
@@
In assessing data fidelity in the image domain, we introduce the residual-to-dirty ratio (RDR). Consider the residual dirty image cube, defined as ${\mathbf{X}}^{\mathrm{res}}=\mathbf{X}^{\mathrm{dirty}}-\mathrm{Re}\{\Phi^{\dagger}(\Phi({\mathbf{X}}))\}$, the overall and channel-wise RDR values can be calculated as
@@
Finally, the spectral index map from a given image cube $\mathbf{X}$, is obtained via linear regression applied at each pixel position to estimate the slope between the logarithm of spectral frequencies and the pixel intensities along the spectral axis following (23). To ensure numerical stability, the spectral index is only evaluated at pixels whose values above heuristic noise levels (14) across all the channels. Consider the groundtruth spectral index map $\overline{\bm{\alpha}}$, the accuracy of the reconstructed spectral index map $\bm{\alpha}$ is measured by the spectral-index SNR metric (sSNR), defined as $\mathrm{SNR(\bm{\alpha,\overline{\bm{\alpha}}})}$.
5.4 Reconstruction results
Channel 1 Image Channel 1 Dirty/Residual Channel 36 Image Channel 35 Dirty/Residual Spectral Index Map
@@ 5.4 Reconstruction results
SNR, logSNR, RDR (2.23)
Refer to captionSNR, logSNR Refer to captionRDR (1.18) Refer to captionSNR, logSNR Refer to captionRDR (5.25) Refer to captionsSNR
@@
23.44, 27.32, 1.91
Refer to caption20.25, 25.46 Refer to caption0.92 Refer to caption30.52, 27.53 Refer to caption4.85 Refer to caption17.19
@@
26.54, 30.16, 2.39
Refer to caption22.09, 30.63 Refer to caption1.10 Refer to caption33.13, 28.48 Refer to caption6.03 Refer to caption16.81
@@
11.84, 18.92, 2.21
Refer to caption8.85, 13.21 Refer to caption1.53 Refer to caption13.21, 12.10 Refer to caption5.01 Refer to caption14.15
@@
25.15, 29.68, 1.97
Refer to caption22.32, 28.67 Refer to caption1.05 Refer to caption30.56, 29.38 Refer to caption4.73 Refer to caption18.65
@@
29.50, 32.84, 2.19
Refer to caption28.25, 31.93 Refer to caption1.06 Refer to caption31.22, 31.55 Refer to caption5.19 Refer to caption18.82
@@
Figure 5: Reconstruction results of a simulated inverse problem using radio galaxy image of 3C 353. Row 1 shows the ground truth. Row 2-6 show results from uSARA, AIRI, WSClean, Hyper-uSARA, and HyperAIRI. Columns 1 and 3 display estimated images at the first and last channels in a logarithmic scale. Columns 2 and 4 display the back-projected dirty images (row 1, normalized to $[0,1]$) or the back-projected residual images (rows 2-6). column 5 illustrates the spectral index maps. Two zoomed-in regions, one focusing on the central black hole and the other on a hotspot to the right, highlight visual differences. Channel-wise SNR and logSNR values (dB) are indicated at at the bottom-right corners of panels in columns 1 and 3, RDR values ($10^{-3}$) in columns 2 and 4 (with ground-truth values in row 1), and sSNR values (dB) in columns 5. Overall metrics for the entire image cube are listed next to each algorithm name, with units matching with the channel-wise metrics. The ground-truth overall RDR value is included in the top-left bracket of row 1.
Firstly, we discuss the overall numerical performance of the various methods on the simulated test dataset, as reported in Table 1, where metrics are computed over the entire image cubes. We also examine channel-wise metrics illustrated in Fig. 4, where each point represents the average across the test cases, and the shaded region denotes $95\%$ confidence interval.
To analyse the fidelity of the reconstructed image cubes to the ground truth, we first examine the SNR metric quantifying how faithfully bright features are recovered. The hyperspectral algorithms Hyper-uSARA and HyperAIRI improve the mean SNR over their respective monochromatic counterparts uSARA and AIRI by about 1 and 1.5 dB. As expected, WSClean’s reconstruction quality ranks lowest, a consequence of smoothing with a channel-specific CLEAN beam and adding with the residual dirty image. Fig. 4 shows that, generally, channel-wise SNR obtained by all algorithms increases with the frequency. This behaviour could be explained by the nature of the hyperspectral Fourier sampling, where larger spatial bandwidths are probed at the highest frequencies. HyperAIRI exploits inter-channel correlations effectively, drastically narrowing the gap in SNR among the low and high frequencies. Similar behaviour is observed with Hyper-uSARA although to a lesser extent. We then examine the logSNR metric results, assessing the recovery of faint emissions, shows that Hyper-uSARA and HyperAIRI exceed their monochromatic counterparts in overall average logSNR by 1.1 and 2.4 dB respectively, with WSClean again lagging behind in quality. The channel-wise logSNR curves in Fig. 4 highlights the overall superior performance of HyperAIRI. Furthermore, we notice a consistent performance of the PnP approaches across channels, with a minor drop in the logSNR values at both ends, whereas the optimisation-based algorithms exhibit logSNR values increasing with the frequency. In contrast, WSClean displays more pronounced drops in logSNR at the lowest and highest frequencies.
@@
Secondly, we conduct a qualitative analysis, by analysing the reconstructed images of a selected inverse problem, whose ground-truth cube is derived from the image of the radio galaxy 3C 353. Fig. 5 shows the reconstructions of the first and last channels obtained with different algorithms. The visual results are generally consistent with the quantitative analysis. By design, WSClean exhibits lower resolution and dynamic range. For the optimisation-based algorithms, Hyper-uSARA exhibits fewer wavelet-like artefacts around the edges of bright structures than its monochromatic counterpart (see the two zoom regions), particularly in the first channel. As for the PnP algorithms, they both exhibit fewer ringing artefacts than uSARA and Hyper-uSARA, especially around the central black hole and the bright emission in the top-right zoom. Furthermore, HyperAIRI produces sharper depiction of relatively bright structures. For instance, the central point source is more compact in the HyperAIRI reconstruction, with a more homogeneous radius across spectral channels. However, compared to Hyper-uSARA, HyperAIRI appears to capture a smooth depiction of the faint extended emission whose intensity values are 3 to 4 orders of magnitude lower than the peak pixel value. This trend, particularly noticeable at the highest frequency, is also reflected in the channel-wise logSNR metric. This could be due to the effective noise level and the resulting choice of the denoiser. Lowering the heuristic noise level in the highest-frequency channel may help restore those faint features in this case.
@@
The spectral index maps produced by all methods generally agree with the ground truth. Upon closer inspection, the spectral index maps of WSClean exhibit fake details which strongly correlate with image spatial features (see the two zooms). This effect arises from CLEAN’s high sensitivity to the varying spatial resolution across channels. In contrast, the spectral index map of both Hyper-uSARA and HyperAIRI are more faithful to the ground truth, with both algorithms substantially reducing the artefacts present in their monochromatic counterparts. When comparing these two hyperspectral algorithms, the spectral index maps of HyperAIRI shows even milder ringing artefacts (see the zooms of the central black hole and the right hotspot) but appears marginally more grainy in the bottom region, which is linked to the smoothness of its reconstruction in the higher-frequency channels.
The computational details of various algorithms are indicated in Table 1. As expected, WSClean appears to be 5 and 4 times faster than optimisation algorithms and PnP algorithms respectively. Its computational efficiency is explained by its small number of passes through the measurement operators within its major cycle. In comparison, the rest of the algorithms, sharing the same iterative structure, are highly iterative in nature. On average, both optimisation-based algorithms require comparable computational time, despite Hyper-uSARA demanding nearly $50\%$ more iterations than uSARA. This highlights the efficiency of the hyperspectral regularisation, whose sub-iterative proximity operator converges at least twice as fast. PnP algorithms also exhibit comparable computational time. Although HyperAIRI denoisers are more complex than AIRI, they do not introduce additional computational cost. Notably, PnP approaches generally require more iterations than pure optimisation methods. Nevertheless, they remain faster overall due to the efficient GPU-based inference of their DNN denoisers. In addition, the gradient steps of the PnP algorithms are slightly faster than those of the optimisation-based algorithms, primarily because of differences in CPU performance and memory configuration between the CPU and GPU nodes on Cirrus.

### v2

@@ V.1 Test dataset
The considered ground-truth images are derived from four radio images of size $N=512\times 512$, which consist of giant radio galaxies 3C 353 (M. R. Swain et al., 1998) and Messier 106 (T. Shimwell et al., 2022), and the radio galaxy clusters Abell 2034 and PSZ2 G165.68+44.01 (A. Botteon et al., 2022) shown in Figure 3. The images are first preprocessed to reduce noise and artifacts, then randomly exponentiated to make the final dynamic range in the interval $10^{3},10^{5}]$. We follow the same preprocessing strategies for OAID and MRID and create 40 HDR monochromatic images. Following P.-A. Thouvenin et al. (2023a), for each image, a spectral index map and a curvature map are generated, in a manner consistent with the procedure used for the training dataset. Specifically, the spectral index map is modeled as the sum of a blurred version of the ground truth in logarithmic scale and a Gaussian field, clipped to the range [5,0][-5,0]. The curvature map is modeled as a Gaussian field, clipped to the range [0.5,0.5][-0.5,0.5]. Each image cube contains $L=36$ channels, with the first channel fixed at the observation frequency of 1 GHz and the last channel randomly chosen from the interval $[1.5,2.0]$ GHz. The observation frequencies of the remaining channels are uniformly spaced within this range. We simulate 40 ground-truth image cubes of size $512\times 512\times 36$ following the spectral model (1).
@@
Hyperspectral RI data are simulated using the antenna configurations of the VLA. The telescope is known for its unique ability to change the positions of the antennas along rail tracks, forming four standard configurations with maximum baselines ranging from 36.4 km to 1.03 km. Configurations with larger baselines provide higher angular resolution, while those with shorter baselines offer higher sensitivity to spatially extended emission. Astronomers often combine measurements from different configurations to obtain a comprehensive Fourier sampling of the target object. Under these considerations, we simulate RI observations combining configurations A and C of the VLA. The Fourier sampling patterns for both configurations are generated using the MeqTrees software package (J. E. Noordam & O. M. Smirnov, 2010). We randomly vary the telescope’s pointing direction, and the observation duration within the range $[5,10]$ hours, and fix the integration time to 12 seconds. The resulting total number of points in each sampling pattern, at any given spectral channel, varies from $1.1\times 10^{6}$ to $2.1\times 10^{6}$. Examples of simulated sampling patterns for spectral channels with different observation frequencies are shown in Figure 1. The hyperspectral RI data are simulated under the assumption of a narrow field of view, and following (4), where the standard deviation of the observation noise is fixed across the spectral channels aligning with the target dynamic range of the cube as per the heuristic (14), such that the target dynamic range of the reconstructed image cube matches that of the corresponding ground-truth image cube. The observation’s nominal resolution is parameterized by the super-resolution factor, defined as the ratio between the image spatial Fourier bandwidth and the maximum projected baseline. We fix the super-resolution factor at the highest-frequency channel to 1.5, leading to a super-resolution factor across different channels in the range $[1.5,3.0]$. The resulting test dataset is composed of 40 inverse problems.
V.2 Benchmark settings
@@
For benchmarking, we compare HyperAIRI’s performance against four algorithms. These are its optimization-counterpart Hyper-uSARA, the hyperspectral CLEAN variant in WSClean (A. Offringa & O. Smirnov, 2017), and the monochromatic algorithms AIRI and uSARA which are applied for each channel independently. Briggs weighting with robustness parameter set to 0 is considered for all algorithms.
@@
In our experiments, the regularization parameters for the optimization algorithms and the effective noise levels for PnP algorithms are directly linked to the heuristic noise level via (14) without further tuning. WSClean parameters are tuned on a small subset of the test dataset and then fixed across the entire dataset to ensure the best overall image quality. In particular, the auto-masking and auto-thresholding values are set to 1.0 and 0.5, respectively, and the number of polynomial terms for spectrum fitting is fixed at 4. The full WSClean command is shown in Appendix B. We also note that the WSClean reconstructions are normalized by the area of the CLEAN beam $A_{\mathrm{beam}}$ on a per-channel basis, thus converting the unit of pixel intensities from Jy/beam to Jy/pixel to align with the ground-truth image cube. The normalization factor is given by
@@
HyperAIRI, Hyper-uSARA, AIRI and uSARA are . All experiments were run on Cirrus, a UK Tier 2 HPC service that provides access to both CPU and GPU compute nodes. Each CPU node comprises two Intel 18-core Xeon E5-2695 processors with 256 GB of memory, while each GPU node comprises two 20-core Intel Xeon Gold 6148 processors, four NVIDIA Tesla V100-SXM2-16GB GPUs, and 384 GB of memory. For each inverse problem in the test dataset, uSARA, WSClean and Hyper-uSARA run on a single CPU node, whereas AIRI and HyperAIRI were executed on a single GPU node to accelerate the inference of DNNs. Each AIRI and HyperAIRI denoisers is trained with half GPU node.
Algorithm Joint SNR logSNR RDR sSNR Iterations Resources ${t}_{\text{total}}$ ${t}_{\text{gradient}}$ ${t}_{\text{denoising}}$ deconv. $\pm$ std (dB) $\pm$ std (dB) $\pm$ std $(\times 10^{-3})$ $\pm$ std (dB) $\pm$ std CPUs GPUs $\pm$ std (hr) $\pm$ std (s) $\pm$ std (s) uSARA No 28.32 $\pm$ 3.24 21.17 $\pm$ 3.56 2.27 $\pm$ 1.32 9.54 $\pm$ 2.34 1251 $\pm$ 299 36 0 2.60 $\pm$ 0.55 2.10 $\pm$ 0.27 5.24 $\pm$ 6.73 AIRI No 28.70 $\pm$ 2.36 20.67 $\pm$ 5.01 2.62 $\pm$ 1.41 9.18 $\pm$ 3.90 3000 $\pm$ 0 36 4 1.84 $\pm$ 0.14 1.60 $\pm$ 0.23 0.46 $\pm$ 0.32 WSClean Yes 11.27 $\pm$ 1.57 12.05 $\pm$ 4.08 2.36 $\pm$ 1.27 3.81 $\pm$ 5.91 10 $\pm$ 1 36 0 0.56 $\pm$ 0.24 Hyper-uSARA Yes 29.30 $\pm$ 2.91 22.30 $\pm$ 4.15 2.32 $\pm$ 1.34 11.35 $\pm$ 3.78 1827 $\pm$ 675 36 0 2.42 $\pm$ 0.87 2.09 $\pm$ 0.27 2.58 $\pm$ 1.43 HyperAIRI Yes 30.53 $\pm$ 1.87 23.07 $\pm$ 5.13 2.44 $\pm$ 1.36 10.87 $\pm$ 3.91 3000 $\pm$ 0 36 4 1.90 $\pm$ 0.14 1.61 $\pm$ 0.23 0.52 $\pm$ 0.25
V.3 Evaluation metrics
@@
To evaluate the algorithms’ performance in reconstructing faint emissions, we evaluate the SNR on logarithmic scale images, denoted as logSNR. Given a ground-truth image cube, with a dynamic range $a$, we apply the logarithmic scale transform parameterized by $a$ such that
@@
In assessing data fidelity in the image domain, we introduce the residual-to-dirty ratio (RDR). overall and channel-wise RDR values can be calculated as
@@
Finally, the spectral index map from a given image cube $\mathbf{X}$, is obtained via linear regression applied at each pixel position to estimate the slope between the logarithm of spectral frequencies and the pixel intensities along the spectral axis following (23). To ensure numerical stability, the spectral index is only evaluated at pixels whose values above heuristic noise levels (14) across all the channels. Consider the ground-truth spectral index map $\overline{\bm{\alpha}}$, the accuracy of the reconstructed spectral index map $\bm{\alpha}$ is measured by the spectral-index SNR metric (sSNR), defined as $\mathrm{SNR(\bm{\alpha,\overline{\bm{\alpha}}})}$.
V.4 Reconstruction results
Channel 1 Channel 36 Spectral
Image Dirty/Residual Image Dirty/Residual Index Map
@@ V.4 Reconstruction results
[…]
@@
Refer to caption20.25, 25.46 $0.92\!\times\!10^{-3} Refer to caption$30.52, 27.53 $4.85\!\times\!10^{-3} Refer to caption$17.19
@@
Refer to caption22.09, 30.63 $1.10\!\times\!10^{-3} Refer to caption$33.13, 28.48 $6.03\!\times\!10^{-3} Refer to caption$16.81
@@
Refer to caption8.85, 13.21 $1.53\!\times\!10^{-3} Refer to caption$13.21, 12.10 $5.01\!\times\!10^{-3} Refer to caption$14.15
@@
Refer to caption22.32, 28.67 $1.05\!\times\!10^{-3} Refer to caption$30.56, 29.38 $4.73\!\times\!10^{-3} Refer to caption$18.65
@@
Refer to caption28.25, 31.93 $1.06\!\times\!10^{-3} Refer to caption$31.22, 31.55 $5.19\!\times\!10^{-3} Refer to caption$18.82
@@
Figure 5: Reconstruction results of a simulated inverse problem using the image of 3C 353. Row 1 shows the ground-truth images of channels 1 and 36 (logarithmic scale; columns 1, 3), the associated dirty images (linear scale; columns 2, 4), and the ground-truth spectral index map (linear scale; column 5). Rows 2–6 present the results of all algorithms: estimated images of the selected channels (logarithmic scale; columns 1, 3), associated residual dirty images (linear scale; columns 2, 4), and estimated spectral index map (linear scale; column 5). Two zoomed-in regions are provided, highlighting (i) the central black hole and (ii) the eastern hotspot. Quantitative metrics are indicated at the bottom-right of the corresponding panels.
Firstly, we discuss the overall numerical performance of the various methods on the simulated test dataset, as reported in Table 1, where metrics are computed over the entire image cubes. We also examine channel-wise metrics illustrated in Figure 4, where each point represents the average across the test cases, and the shaded region denotes $95\%$ confidence interval.
To analyze the fidelity of the reconstructed image cubes to the ground truth, we first examine the SNR metric quantifying how faithfully bright features are recovered. The hyperspectral algorithms Hyper-uSARA and HyperAIRI improve the mean SNR over their respective monochromatic counterparts uSARA and AIRI by about 1 and 1.5 dB. As expected, WSClean’s reconstruction quality ranks lowest, a consequence of smoothing with a channel-specific CLEAN beam and adding with the residual dirty image. Figure 4 shows that, generally, channel-wise SNR obtained by all algorithms increases with the frequency. This behavior could be explained by the nature of the hyperspectral Fourier sampling, where larger spatial bandwidths are probed at the highest frequencies. HyperAIRI exploits inter-channel correlations effectively, drastically narrowing the gap in SNR among the low and high frequencies. Similar behavior is observed with Hyper-uSARA although to a lesser extent. We then examine the logSNR metric results, assessing the recovery of faint emissions, shows that Hyper-uSARA and HyperAIRI exceed their monochromatic counterparts in overall average logSNR by 1.1 and 2.4 dB respectively, with WSClean again lagging behind in quality. The channel-wise logSNR curves in Figure 4 highlights the overall superior performance of HyperAIRI. Furthermore, we notice a consistent performance of the PnP approaches across channels, with a minor drop in the logSNR values at both ends, whereas the optimization-based algorithms exhibit logSNR values increasing with the frequency. In contrast, WSClean displays more pronounced drops in logSNR at the lowest and highest frequencies.
@@
Secondly, we conduct a qualitative analysis, by analyzing the reconstructed images of a selected inverse problem, whose ground-truth cube is derived from the image of the radio galaxy 3C 353. Figure 5 shows the reconstructions of the first and last channels obtained with different algorithms. The visual results are generally consistent with the quantitative analysis. By design, WSClean exhibits lower resolution and dynamic range. For the optimization-based algorithms, Hyper-uSARA exhibits fewer wavelet-like artifacts around the edges of bright structures than its monochromatic counterpart (see the two zoom regions), particularly in the first channel. As for the PnP algorithms, they both exhibit fewer ringing artifacts than uSARA and Hyper-uSARA, especially around the central black hole and the bright emission in the top-right zoom. Furthermore, HyperAIRI produces sharper depiction of relatively bright structures. For instance, the central point source is more compact in the HyperAIRI reconstruction, with a more homogeneous radius across spectral channels. However, compared to Hyper-uSARA, HyperAIRI appears to capture a smooth depiction of the faint extended emission whose intensity values are 3 to 4 orders of magnitude lower than the peak pixel value. This trend, particularly noticeable at the highest frequency, is also reflected in the channel-wise logSNR metric. This could be due to the effective noise level and the resulting choice of the denoiser. Lowering the heuristic noise level in the highest-frequency channel may help restore those faint features in this case.
@@
The spectral index maps produced by all methods generally agree with the ground truth. Upon closer inspection, the spectral index maps of WSClean exhibit fake details which strongly correlate with image spatial features (see the two zooms). This effect arises from CLEAN’s high sensitivity to the varying spatial resolution across channels. In contrast, the spectral index map of both Hyper-uSARA and HyperAIRI are more faithful to the ground truth, with both algorithms substantially reducing the artifacts present in their monochromatic counterparts. When comparing these two hyperspectral algorithms, the spectral index maps of HyperAIRI shows even milder ringing artifacts (see the zooms of the central black hole and the right hotspot) but appears marginally more grainy in the bottom region, which is linked to the smoothness of its reconstruction in the higher-frequency channels.
The computational details of various algorithms are indicated in Table 1. As expected, WSClean appears to be 5 and 4 times faster than optimization algorithms and PnP algorithms respectively. Its computational efficiency is explained by its small number of passes through the measurement operators within its major cycle. In comparison, the rest of the algorithms, sharing the same iterative structure, are highly iterative in nature. On average, both optimization-based algorithms require comparable computational time, despite Hyper-uSARA demanding nearly $50\%$ more iterations than uSARA. This highlights the efficiency of the hyperspectral regularization, whose sub-iterative proximity operator converges at least twice as fast. PnP algorithms also exhibit comparable computational time. Although HyperAIRI denoisers are more complex than AIRI, they do not introduce additional computational cost. Notably, PnP approaches generally require more iterations than pure optimization methods. Nevertheless, they remain faster overall due to the efficient GPU-based inference of their DNN denoisers. In addition, the gradient steps of the PnP algorithms are slightly faster than those of the optimization-based algorithms, primarily because of differences in CPU performance and memory configuration between the CPU and GPU nodes on Cirrus.
VI Validation on real ASKAP observations
In this section, we revisit the ASKAP measurements of the field SB9442-35 previously imaged by A. G. Wilber et al. (2023a, b) with uSARA and AIRI. To evaluate the performance of proposed hyperspectral algorithms, Hyper-uSARA and HyperAIRI, we support our analysis with the reconstructions obtained by their monochromatic counterparts, as well as the hyperspectral CLEAN variant in WSClean.
VI.1 Data details for the field SB9442-35